THERMOCHEMICAL INFORMATION AND ROTATIONAL STRUCTURE IN THE AUTOIONIZATION SPECTRA OF HCO.

\underline{ERIC J. ZÜCKERMAN}, ROBERT J. FOLTYNOWICZ, MICHAEL KONOPKA, HARTMUT G. HEDDERICH and EDWARD R. GRANT, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393; SHIH-HUI JEN and I-CHIA CHEN, Department of Chemistry, National Tsing Hua University, Hsinchu, Taiwan, 30043, Republic of China.

Rotationally resolved autoionization spectra of the HCO radical in high-Rydberg states built on the (010) core vibrational state have been obtained from selected rovibrational levels of the 3pπ^{2}Π (010) Σ^{+} Rydberg state. Simulations of the spectra lead to reliable ionization potentials for detailed rovibrational states of HCO^{+}. Data for bending overtones supply information on anharmonic terms in the cation vibrational potential. Rotational structure associated with autoionizing states is analyzed.